[Brahms-l] Shift report 20040513 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Thu May 13 2004 - 08:12:11 EDT
------------------------------------------------------------
Date: Thu May 13 08:12:08 2004, Shift: 00:00-08:00
Supervisor: hiro, with: 
Log book#15 pages 120 - 120

Shift Summary:
------------------------------------------------------------
1. beam for 6hr
   FS
   1/1 B 2M already completed.
   1/2 B 0.2M already completed.
   0 field already completed.
   1/1 A 0.5M completed.
   We decided to take more 1/1A to match 1/1B.



2. 11856-11859
3. T3 A3-xu does not hold HV.
   D5 set value and measured value is different. 


------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu May 13 08:12:52 2004

This archive was generated by hypermail 2.1.8 : Thu May 13 2004 - 08:13:17 EDT